OpenStudy (anonymous):

80 percent of the 160 4th grade girls had long hair, and 85 percent of the 120 5th grade girls had long hair. Which class had more girls with long hair, and by how many?

OpenStudy (anonymous):

. 4th grade, 5 more B. 4th grade, 26 more C. 5th grade, 5 more D. 5th grade, 26 more her are your choses

OpenStudy (anonymous):

i mean here

OpenStudy (anonymous):

i think it is A

OpenStudy (anonymous):

80% of 160 is 128 85% of 120 is 102 So 128 - 102 = 26 Obviously there are 26 girls more in 4th grade with long hair... Fair enough? :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Do you know how to get 80% of 160?

OpenStudy (anonymous):

no i do not

OpenStudy (anonymous):

Ok, you know that there are always 1-100% So ANY number you are asked, just divide it but 100 first, and you will get value for 1% And then multiply it by number you need, in this case 80... 160/100*80 = 128 120/100*85 = 102 If you need 5% 160/100*5 If you need 0.5% 160/100*0.5 Ok? :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

Ok, good luck! :)

OpenStudy (anonymous):

thanks bye